# When Collaborative Maintenance Falls Short: The Persistence of Retracted Papers on Wikipedia

### v2

@@ 3.6.2 Retraction Event Features
This feature group captures characteristics of the retraction event itself. These features provide context on the severity of the paper’s issues and the timeline of its withdrawal, which may signal to editors the urgency of a correction and thus influence its speed. Data for these features were retrieved from OpenAlex priem2022openalex and Retraction Watch oransky2012retraction .
  • Time to Retraction. Time to retraction is calculated as the time between a paper’s publication date and its official retraction date, retrieved from OpenAlex and Retraction Watch. Longer times between publication and retraction may delay the correction of citations on Wikipedia, as outdated citations may become embedded in the platform’s content. Prior studies have highlighted the negative impact of delayed retractions on scientific credibility and public trust brainard2018massive .
@@
  • Retraction Reason. According to the framework presented in Zhang et al., the retraction reason can be categorized into 8 groups: “plagiarism”, “self-plagiarism”, “falsification/manipulation”, “error & unreliable results”, “authorship issues”, “ethical issues”, “others”, and “not available/lack of information” zhang2020collaboration . For this study, we focused specifically on falsification/manipulation (fraud) and error & unreliable results (error). Following this framework, we searched for the keywords “falsification,” “manipulation,” and “misconduct” in the Retraction Watch reasons to identify fraud papers, and we used the keywords “error,” “contamination,” and “unreliable” to identify error papers. These fraud and error categories were then combined into a broader category labeled “fraud & errors,” capturing retractions that may draw more editorial attention. Papers falling into this category were assigned a binary value of 1.
@@ 3.6.3 Wikipedia Page Features
This feature group captures characteristics of Wikipedia pages that may influence the speed with which retracted citations are identified and corrected. We aim to understand how Wikipedia’s collaborative and dynamic editorial processes impact the persistence of problematic citations. All features in this group are derived from Wikimedia data, retrieved through the Wikimedia REST MediaWiki and MediaWiki Action WikiAction APIs.
  • Number of Unique Editors. This variable serves as a proxy for the level of collaborative oversight on a given Wikipedia page. To ensure this was a valid baseline covariate and to avoid look-ahead bias, we measured it dynamically. For each citation pair, we calculated the number of distinct human editors who had contributed to the page up to the start of that citation’s risk period (i.e., the point after the cited paper was both retracted and present on the page). To ensure the measure reflects meaningful editorial engagement, we excluded bot and anonymous edits from this count. The rationale is that a page with a larger community of editors at the time a citation becomes at risk may experience greater collaborative oversight, potentially leading to a faster correction.
  • Controversy Noted at Initial Citation. This feature indicates whether the Wikipedia article describes the cited paper as controversial or disputed at the time it was first added to the page. We identify this by manually checking whether the initial citation frames the paper as debated, flawed, or otherwise questionable. Such editorial framing may signal early skepticism and influence the speed at which the citation is later corrected.
We excluded revision count at risk time from the model due to its near-perfect correlation with the number of unique editors at risk time.
@@ 3.7 Data Preprocessing and Model Specification
Before model fitting, we preprocessed the independent variables to ensure model stability. We applied a log-transformation (log(x+1)) to continuous variables exhibiting high positive skew, including academic citation count, team size, tweet mention count, time to retraction, and number of unique editors. Subsequently, these same continuous variables were standardized using z-score normalization to allow for the comparison of their effect sizes. Binary and dummy variables were left unchanged. Our initial sample of 845 problematic citations was reduced to a final analytical sample of 764 for our models after excluding “proactive corrections”—cases where the correction was made before the official retraction date.
Our analysis proceeds in two stages: first, we assess the suitability of the standard Cox Proportional Hazards model (PH), and second, based on diagnostic tests, we select and specify a more appropriate Accelerated Failure Time (AFT) model.
3.7.1 Model Selection
We began by fitting a Cox PH model, the semi-parametric standard for survival analysis. We then tested the model’s core assumption of proportional hazards using a statistical test based on scaled Schoenfeld residuals schoenfeld1982partial . The test revealed that our categorical variable for the academic domain exhibited non-proportional hazards (see Appendix Table A1). Specifically, the dummy variable representing the Life Sciences domain failed the test ($p=0.0347$), indicating that its effect on the hazard of correction is not constant over time.
Given this violation of a core assumption, we concluded that the Cox PH model was not the most appropriate choice for our data. We therefore turned to the family of parametric Accelerated Failure Time (AFT) models. Unlike the Cox model, which models the hazard rate, AFT models directly model the time to an event and do not rely on the proportional hazards assumption, making them a robust alternative in this context.
3.7.2 AFT Model Specification
Let $T_{i}$ denote the time (in days) between a citation $i$ entering its risk period to the first correction event on Wikipedia. To model this, we specified a general Accelerated Failure Time (AFT) model, which links the logarithm of the survival time to a set of covariates:
$\log T_{i}=\beta_{0}\;$ $+\beta_{1}\,\mathrm{OpenAccess}_{i}+\beta_{2}\,\mathrm{CitationsAtRisk}_{i}+\beta_{3}\,\mathrm{TeamSize}_{i}$
$+\beta_{4}\,\mathrm{PhysSci}_{i}+\beta_{5}\,\mathrm{LifeSci}_{i}+\beta_{6}\,\mathrm{SocSci}_{i}+\beta_{7}\,\mathrm{TweetMentions}_{i}$
$+\beta_{8}\,\mathrm{TimeToRetraction}_{i}+\beta_{9}\,\mathrm{PreRetractionEditorial}_{i}+\beta_{10}\,\mathrm{RetractionReason}_{i}$
$+\beta_{11}\,\mathrm{UniqueEditors}_{i}+\beta_{12}\,\mathrm{ControversyNoted}_{i}+\varepsilon_{i},$ (1)
@@ 3.7.2 AFT Model Specification
  • $\textit{OpenAccess}_{i}$: indicator if the paper was published open access;
  • $\textit{CitationsAtRisk}_{i}$: academic citation count of the paper at the start of the risk period;
  • $\textit{TeamSize}_{i}$: number of authors of the retracted paper;
  • $\textit{PhysSci}_{i}$, $\textit{LifeSci}_{i}$, $\textit{SocSci}_{i}$: academic domain dummies (Health Sciences as reference);
  • $\textit{TweetMentions}_{i}$: number of Twitter/X mentions at the start of the risk period;
  • $\textit{TimeToRetraction}_{i}$: days from publication to official retraction;
  • $\textit{PreRetractionEditorial}_{i}$: indicator for any editorial action prior to retraction (e.g., expression of concern);
  • $\textit{RetractionReason}_{i}$: salient retraction reason (fraud/manipulation or error/unreliable results);
  • $\textit{UniqueEditors}_{i}$: number of distinct human editors contributing before the risk period;
  • $\textit{ControversyNoted}_{i}$: indicator if the citation was framed as controversial at the time of addition;
  • $\varepsilon_{i}$: an error term following an assumed parametric distribution.
To select the most appropriate parametric distribution for our data, we fitted three distinct AFT models—Weibull, Log-Normal, and Log-Logistic—which are distinguished by the assumed distribution of the error term $\varepsilon_{i}$. We then compared their relative goodness of fit using the Akaike Information Criterion (AIC), a standard metric for model selection that penalizes model complexity akaike2003new . The Weibull AFT model yielded the lowest AIC and was therefore selected for our final analysis (see Appendix Table A2). All survival analyses were conducted in Python using the lifelines package davidson2019lifelines .
@@ 4.1 Overview of Retracted Papers on Wikipedia
Figure 3: Share of problematic (orange) vs unproblematic (green) citations on Wikipedia.
@@ 4.1.1 Problematic Citations
We analyze the distribution of different types of retracted paper citations on Wikipedia. As shown in Figure 3, among the 1,181 citation pairs found, 845 were identified as problematic citations, with 608 cited before retraction and 237 cited after retraction without proper in-text warnings. This means that 71.6% of all collected citations represent problematic cases, where retracted papers were either cited before their retraction status became known or cited afterwards without adequate warnings to readers. 51.5% of all citations fall into the former category. These citations often persist due to a lack of systematic monitoring, leaving articles exposed to outdated or invalidated information. In contrast, 28.3% of the citations were non-problematic, meaning that their retracted status was appropriately mentioned at the time of citation. Critically, the category of citations introduced after a paper’s retraction but without proper in-text warnings constitutes 20.2% of all citations in our sample, revealing the limitations of current editorial and automated interventions in flagging and fixing such citations.
@@ 4.1.2 Citation of Retracted vs Non-retracted Papers Across Academic Domains
The Health Sciences domain stands out with the highest ratio of retracted papers cited on Wikipedia, at 0.0220, compared to a much lower 0.0064 ratio for non-retracted papers. This disparity suggests a heightened vulnerability of this domain to problematic citations. Since retracted Health Sciences papers are 3.44 times more likely to be cited on Wikipedia than non-retracted ones, formal retractions in this high-stakes field are not reliably reflected in Wikipedia’s citation practices.
@@ 4.2.1 Distribution of the Corrections
Figure 6: Kaplan-Meier survival curve of problematic citations. The y-axis shows the proportion of citations that remain uncorrected over time. The x-axis represents the time to correction in days, measured from the start of the risk period for each citation, which is defined as the later of the paper’s official retraction date or the citation’s addition to Wikipedia.
@@ 4.2.2 Descriptive Analysis of Correction Times
Our final analytical sample consists of 764 problematic citations. Of these, 486 were corrected during our observation period (the event), while 278 remained uncorrected and were right-censored at the end of the study. To visualize this process, we use the Kaplan-Meier estimator kaplan1958nonparametric , a standard non-parametric method for plotting the probability of an event not yet occurring over time while properly accounting for censored data. The resulting survival curve for our sample is shown in Figure 6.
@@
Figure 7: Coefficients from the Weibull Accelerated Failure Time (AFT) model predicting the time to correction for problematic citations (Concordance = 0.671; log-likelihood ratio test: $p<0.001$). Feature categories group coefficients. Bars indicate the coefficient estimates; error bars represent a $95\%$ confidence interval. The shapes and colors of the data points denote feature groups. Features with negative coefficients are associated with an accelerated time to correction (i.e., faster correction), while positive coefficients indicate a decelerated time (i.e., slower correction).
4.2.3 Factors Influencing the Time to Correction: AFT Model Results
To identify factors that predict the time until a problematic citation is corrected, we fitted an AFT model with an assumed Weibull distribution for the survival times, which directly models the speed of correction. The results are interpreted as time ratios, where a ratio below 1 indicates a faster correction and a ratio above 1 indicates a slower correction.
The overall model is statistically significant ($\text{log-likelihood ratio test}=117.551$, $p<0.001$) and demonstrates moderate predictive accuracy with a concordance index of 0.671. While some independent variables are modestly correlated (see Appendix Figure B1), all variance inflation factors (VIFs) are well below the standard threshold of 5 (see Appendix Table B1), indicating that multicollinearity is not a concern. The following sections detail the effects of specific paper-, retraction event-, and Wikipedia page-level features on the time to correction. Figure 7 presents the results, which are discussed below, broken down by feature groups.

### v3

@@ 3.6.2 Retraction Event Features
This feature group captures characteristics of the retraction event itself. These features provide context on the severity of the paper’s issues and the timeline of its withdrawal, which may be associated with how quickly citations are corrected. Data for these features were retrieved from OpenAlex Priem et al. (2022) and Retraction Watch Oransky and Marcus (2012).
  • Time to Retraction. Time to retraction is calculated as the time between a paper’s publication date and its official retraction date, retrieved from OpenAlex and Retraction Watch. Longer times between publication and retraction may be associated with slower correction, as outdated citations may become embedded in the platform’s content. Prior studies have highlighted the negative impact of delayed retractions on scientific credibility and public trust Brainard et al. (2018).
@@
  • Retraction Reason. According to the framework presented in Zhang et al. (2020), the retraction reason can be categorized into 8 groups: “plagiarism”, “self-plagiarism”, “falsification/manipulation”, “error & unreliable results”, “authorship issues”, “ethical issues”, “others”, and “not available/lack of information”. For this study, we focused specifically on falsification/manipulation (fraud) and error & unreliable results (error). We prioritized these categories because they most directly undermine the epistemic validity of the scientific claims. In contrast, retractions due to authorship problems, ethical issues, or other administrative errors, while serious violations of academic norms, do not necessarily imply that the scientific findings themselves are factually incorrect. Following this framework, we searched for the keywords “falsification,” “manipulation,” and “misconduct” in the Retraction Watch reasons to identify fraud papers, and we used the keywords “error,” “contamination,” and “unreliable” to identify error papers. These fraud and error categories were then combined into a broader category labeled “fraud & errors,” capturing retractions that may draw more editorial attention. Papers falling into this category were assigned a binary value of 1.
@@ 3.6.3 Wikipedia Page Features
Following prior work on language-agnostic Wikipedia article features Das et al. (2024), we include a set of page-level characteristics measured at the start of each citation’s risk period. This feature group captures characteristics of Wikipedia pages associated with variation in the timing of reader-facing corrections. We aim to understand how Wikipedia’s collaborative and dynamic editorial processes are associated with the persistence of problematic citations. All features in this group are derived from Wikimedia data, retrieved through the Wikimedia REST MeidaWiki (2025) and MediaWiki Action MediaWiki (2024) APIs.
  • RetractionBot Intervention. We treat the presence of a RetractionBot warning as a time-dependent covariate. We recorded the exact timestamp when the bot first flagged the citation pair. In our time-varying covariate construction, the status is 0 before this timestamp and 1 after for the citation pair. This specification allows us to estimate how the hazard of correction differs before versus after a RetractionBot flag.
  • Number of Unique Editors. This variable serves as a proxy for the level of collaborative oversight on a given Wikipedia page. To ensure this was a valid baseline covariate and to avoid look-ahead bias, we measured it dynamically. For each citation pair, we calculated the number of distinct human editors who had contributed to the page up to the start of that citation’s risk period (i.e., the point after the cited paper was both retracted and present on the page). To ensure the measure reflects meaningful editorial engagement, we excluded bot and anonymous edits from this count. The rationale is that a page with a larger community of editors at the time a citation becomes at risk may experience greater collaborative oversight, potentially correlating with faster correction.
  • Controversy Noted at Initial Citation. This feature indicates whether the Wikipedia article describes the cited paper as controversial or disputed at the time it was first added to the page. We identify this by manually checking whether the initial citation frames the paper as debated, flawed, or otherwise questionable. Such editorial framing may capture early skepticism and be associated with earlier correction.
  • Reference Count. We use the number of references on the Wikipedia page as a proxy for sourcing density and page complexity. To avoid look-ahead bias, we measure this feature dynamically at the start of each citation’s risk period. Pages with more references may reflect stronger sourcing norms and editorial attention, but they may also impose a higher maintenance burden, potentially slowing the identification and correction of problematic citations.
  • Categories Count. This feature captures how broadly the Wikipedia page is situated within the platform’s topical classification system, measured as the number of categories assigned to the page at the start of the citation’s risk period. A larger number of categories may increase the likelihood that the page is monitored by multiple interest groups or WikiProjects, potentially correlating with earlier detection and correction.
  • Media Count. We measure the number of media items (e.g., images and other embedded files) included on the page at the start of the citation’s risk period. Media-rich pages may attract more readership and editorial attention, which may be associated with earlier correction, although they may also be longer or more complex.
We excluded revision count at risk time from the model because it was nearly perfectly correlated with the number of unique editors at risk time. We omit other size proxies (e.g., page length, headings, and wikilinks) because they are highly collinear with these structural measures and would add redundancy.
@@ 3.7 Data Preprocessing and Model Specification
Before model fitting, we preprocessed the independent variables to ensure model stability. We applied a log-transformation $\log(x+1)$ to right-skewed count-like variables, including academic citation count, team size, tweet mention count, time to retraction, and number of unique editors. Subsequently, these continuous variables were standardized using z-score normalization to facilitate comparison of coefficient magnitudes (i.e., log-hazard ratios) across covariates. Binary and dummy variables were left unchanged. Our initial sample of 845 problematic citations was reduced to a final analytical sample of 762 after excluding 83 “proactive corrections”—cases where the correction was made before the official retraction date. We exclude these cases because our risk-set definition begins only once a paper is officially retracted; corrections made prior to retraction likely reflect a different process (e.g., pre-retraction controversy coverage) and would yield negative durations under this operationalization.
[…]
3.7.1 Construction of Time-Dependent Covariate: RetractionBot
A critical methodological challenge in analyzing the association of RetractionBot flagging with correction timing is the temporal nature of its intervention. Treating the bot as a static binary variable would introduce immortal time bias Suissa (2008), as citations must “survive” (remain uncorrected) long enough to be visited by the bot. To rigorously address this, we structured our dataset using the counting process format (start-stop format) Therneau and Grambsch (2000). We split the observation period of any citation targeted by the bot into two distinct intervals:
  • Pre-intervention interval $(0,t_{\mathrm{bot}}]$ where the citation is treated as unexposed ($RetractionBot=0$).
  • Post-intervention interval $(t_{\mathrm{bot}},t_{\mathrm{end}}]$ where it is treated as exposed ($RetractionBot=1$).
Here, $t_{\mathrm{bot}}$ is the timestamp of the first RetractionBot flag for the citation pair, and $t_{\mathrm{end}}$ is the event time (correction) or the censoring time (October 24, 2024), whichever comes first. This transformation allows the risk set to be updated instantaneously at the moment of the bot’s intervention. Citations never visited by the bot contribute a single interval with $RetractionBot=0$ throughout follow-up. After splitting, the analytical dataset contains 1,154 start–stop intervals (rows), representing 762 unique citation pairs, with 484 observed correction events.
3.7.2 Assessment of Proportional Hazards and Model Refinement
We adopted a rigorous, iterative approach to assess the proportional hazards (PH) assumption, ensuring our final model specification accurately reflected the underlying data structure. We assessed the proportional hazards assumption using scaled Schoenfeld residual tests implemented in R (survival::cox.zph), reporting both global and covariate-specific tests.
  • Step 1: Initial Diagnostics. We first fitted a Cox model incorporating the time-dependent RetractionBot intervention and all static variables, using one-hot encoding for Academic Domains (using Health Sciences as reference). Diagnostic tests based on scaled Schoenfeld residuals (see Appendix Table A1) revealed severe violations of the PH assumption. Specifically, the Academic Domain variables ($p<0.001$), Tweet Mention Count ($p<0.001$), Open Access ($p<0.001$), and Retraction Reason (Salient) ($p<0.001$) showed strong deviations. Furthermore, we observed likely spurious violations in several other variables (e.g., Wikipedia Page Reference Count, Categories Count, and Media Count), suggesting that unmodeled heterogeneity in baseline hazards across disciplines may be distorting the model estimates.
  • Step 2: Stratification. To address the violations related to disciplinary differences, we stratified the model by Academic Domain. This modification allows each discipline to have a distinct baseline hazard function while sharing common coefficients for other covariates. Re-evaluation of the PH assumption on this stratified model (see Appendix Table A2) confirmed the effectiveness of this approach: violations for control variables like Media Count ($p=0.569$) and Categories Count ($p=0.1345$) were resolved.
  • Step 3: Time-Dependent Coefficients. However, the stratified diagnostics confirmed that Tweet Mention Count ($p<0.001$), RetractionReason ($p<0.05$), and Open Access ($p<0.05$) continued to significantly violate the PH assumption. This indicates that the effects of social media attention and page sourcing density are not constant but evolve over the risk period. To model these dynamic effects, we specified the final Extended Cox model by including interaction terms for both variables with the logarithm of time ($TweetCount\times\ln t$ and $ReferenceCount\times\ln t$).
3.7.3 Extended Cox Model Specification
The hazard rate for citation $i$ at time $t$ is modeled as:
$h_{i}(t)=h_{0g}(t)\exp\Big($ $\beta_{1}\,\mathrm{RetractionBot}_{i}(t)$
$+\beta_{2}\,\mathrm{TweetCount}_{i}+\gamma_{2}\,\mathrm{TweetCount}_{i}\ln(t+\epsilon)$
$+\beta_{3}\,\mathrm{OpenAccess}_{i}+\gamma_{3}\,\mathrm{OpenAccess}_{i}\ln(t+\epsilon)$
$+\beta_{4}\,\mathrm{TeamSize}_{i}+\beta_{5}\,\mathrm{TimeToRetraction}_{i}+\beta_{6}\,\mathrm{PreRetractionEditorial}_{i}$
$+\beta_{7}\,\mathrm{RetractionReason}_{i}+\gamma_{7}\,\mathrm{RetractionReason}_{i}\ln(t+\epsilon)$
$+\beta_{8}\,\mathrm{CitationsAtRisk}_{i}+\beta_{9}\,\mathrm{ReferenceCount}_{i}+\beta_{10}\,\mathrm{UniqueEditors}_{i}$
$+\beta_{11}\,\mathrm{ControversyNoted}_{i}+\beta_{12}\,\mathrm{MediaCount}_{i}+\beta_{13}\,\mathrm{CategoriesCount}_{i}\Big).$ (1)
@@ 3.7.3 Extended Cox Model Specification
  • $h_{0g}(t)$ is the stratified baseline hazard for academic domain $g$ (Health, Life, Physical, or Social Sciences), allowing baseline risk shapes to vary across domains.
  • $t$ denotes time since the start of the risk period (in days). To avoid $\ln(0)$, we use $\ln(t+\epsilon)$ with $\epsilon=10^{-3}$.
  • $\mathrm{RetractionBot}_{i}(t)$ is a time-dependent covariate indicating whether RetractionBot has become active for citation $i$ by time $t$ (0 before activation, 1 after).
  • $\mathrm{TweetCount}_{i}$ is the tweet mention count at the start of the risk period. Its effect is allowed to vary over time via the interaction $\mathrm{TweetCount}_{i}\times\ln(t+\epsilon)$.
  • $\mathrm{OpenAccess}_{i}$ indicates whether the cited paper is open access. We allow this association to vary over time via $\mathrm{OpenAccess}_{i}\times\ln(t+\epsilon)$.
  • $\mathrm{TeamSize}_{i}$ is the number of authors of the cited paper.
  • $\mathrm{TimeToRetraction}_{i}$ is the number of days from publication to retraction.
  • $\mathrm{PreRetractionEditorial}_{i}$ indicates whether the paper received an editorial notice prior to retraction (e.g., an expression of concern).
  • $\mathrm{RetractionReason}_{i}$ is an indicator for fraud/misconduct/error/unreliable results (1) vs. other reasons (0). Its effect is allowed to vary over time via $\mathrm{RetractionReason}_{i}\times\ln(t+\epsilon)$.
  • $\mathrm{CitationsAtRisk}_{i}$ is the academic citation count at the start of the risk period.
  • $\mathrm{ReferenceCount}_{i}$ is the number of references on the Wikipedia page at the start of the risk period.
  • $\mathrm{UniqueEditors}_{i}$ is the number of unique editors contributing to the Wikipedia page (measured at risk start).
  • $\mathrm{ControversyNoted}_{i}$ indicates whether the citation was initially framed as controversial.
  • $\mathrm{MediaCount}_{i}$ and $\mathrm{CategoriesCount}_{i}$ are Wikipedia page structural features (measured at risk start).
We estimated models using robust (sandwich) standard errors Lin and Wei (1989) clustered by citation pair to account for within-pair dependence introduced by the start–stop (counting-process) representation.
3.7.4 Implementation
We estimated the extended Cox models in R using survival::coxph() with a counting-process (start–stop) specification and robust standard errors clustered at the citation-pair level. Kaplan–Meier curves and survival descriptives were generated and visualized in Python using the lifelines package.
@@ 4.1 Overview of Retracted Papers on Wikipedia
Figure 3: Share of problematic (orange) vs unproblematic (green) retracted-paper citations on Wikipedia.
@@ 4.1.1 Problematic Citations
We analyze the distribution of different types of retracted paper citations on Wikipedia. As shown in Figure 3, among the 1,181 citation pairs found, 845 were identified as problematic citations, with 608 cited before retraction and 237 cited after retraction without proper in-text warnings. This means that 71.6% of all collected citations represent problematic cases, where retracted papers were either cited before their retraction status became known or cited afterwards without adequate warnings to readers. 51.5% of all citations fall into the former category. These citations often persist in the absence of systematic monitoring, leaving articles exposed to outdated or invalidated information. In contrast, 28.4% of the citations were non-problematic, meaning that their retracted status was appropriately mentioned at the time of citation. Critically, the category of citations introduced after a paper’s retraction but without proper in-text warnings constitutes 20.1% of all citations in our sample, revealing the limitations of current editorial and automated interventions in flagging and fixing such citations.
@@ 4.1.2 Citation of Retracted vs Non-retracted Papers Across Academic Domains
The Health Sciences domain stands out with the highest ratio of retracted papers cited on Wikipedia, at 0.0220, compared to a much lower 0.0064 ratio for non-retracted papers. This disparity suggests a heightened vulnerability of this domain to problematic citations. The retracted-to-non-retracted citation ratio is 3.44 times higher in Health Sciences (0.0220 vs. 0.0064), indicating an overrepresentation of retracted work among Wikipedia-cited papers in this domain.
@@ 4.2.1 Distribution of the Corrections
4.2.2 Validation of Time to Correction
To assess the accuracy of our automatically derived correction timestamps, we manually validated the identified correction revisions against human judgments. For in-text retraction acknowledgments ($n=103$), one author validated all cases; for complete removals ($n=464$), we validated a random sample of 100 cases. Concretely, the author opened the revision history for each sampled article–citation pair, located the first revision where the correction is visibly present (either the first stable in-text retraction acknowledgment or the first stable removal of the citation), and compared that revision ID/timestamp to the one returned by our automated pipeline. When the two did not match, the author reviewed adjacent revisions to determine whether the discrepancy was due to rewording of the acknowledgment token/phrase or a switch between correction strategies.
Under our validation criterion—whether the automatically identified correction revision matched the manually identified correction revision (allowing for our three-revision tolerance for brief reversals)—the agreement rate was 87% for in-text acknowledgments and 94% for complete removals. Disagreements primarily arose from editorial rewriting that altered the specific token or phrase used to signal retraction status, or from strategy switches between in-text acknowledgment and complete removal across revisions. We manually corrected these cases in the final analytical dataset used for subsequent analyses.
Figure 6: Kaplan-Meier survival curve of problematic citations. The y-axis shows the proportion of citations that remain uncorrected over time. The x-axis represents the time to correction in days, measured from the start of the risk period for each citation, which is defined as the later of the paper’s official retraction date or the citation’s addition to Wikipedia. Shaded regions represent 95% confidence intervals.
Figure 7: Kaplan–Meier survival curves showing the persistence of problematic retracted paper citations, stratified by academic domain. Curves are shown for descriptive comparison across fields. Domain sample sizes vary, particularly for Social Sciences. Shaded regions represent 95% confidence intervals.
@@ 4.2.3 Descriptive Analysis of Correction Times
Our final analytical sample consists of 762 problematic citations. Of these, 484 were corrected during our observation period (the event), while 278 remained uncorrected and were right-censored at the end of the study. To visualize this process, we use the Kaplan-Meier estimator Kaplan and Meier (1958), a standard non-parametric method for plotting the probability of an event not yet occurring over time while properly accounting for censored data. The resulting survival curve for our sample is shown in Figure 6.
@@
To examine whether this temporal pattern varies across knowledge domains, we next visualize Kaplan-Meier survival curves stratified by academic field. Figure 7 shows clear heterogeneity in correction dynamics across domains.
Citations in the Physical Sciences ($n=100$) exhibit the fastest decline in survival probability, with a median time to correction of approximately 228 days, indicating more rapid correction. In contrast, the Health Sciences ($n=324$) and Life Sciences ($n=278$) display substantially longer persistence and a more pronounced long tail, with median correction times of 1,846 days and 2,720 days, respectively. Social Sciences ($n=60$) show intermediate behavior, with a median of 724 days; greater stepwise variation reflects the smaller sample size. These differences in correction dynamics across academic domains were statistically significant (Log-rank test, $\chi^{2}(3)=77.14,p<0.001$).
Figure 8: Hazard Ratios from the extended Cox Proportional Hazards model predicting the time to correction for problematic citations (Concordance = 0.684; log-likelihood ratio test: $p<0.001$). Feature categories group the predictors. Points indicate the exponentiated coefficient estimates (Hazard Ratios); error bars represent 95% confidence intervals. The shapes and colors distinguish feature groups. Note that for variables exhibiting non-proportional hazards, only the main effect coefficients are visualized here; significant time-dependent interaction terms are detailed in the main text. See Appendix Table A4 for full model results. Hazard Ratios >1.0>1.0 are associated with a higher hazard of correction (i.e., corrected faster), while Hazard Ratios <1.0<1.0 indicate a lower hazard (i.e., the problematic citation persists longer). For continuous predictors, hazard ratios correspond to a one–standard deviation increase in the log-transformed covariate (log1p, then z-scored).
4.2.4 Associations with Time to Correction: Extended Cox Model Results
To identify factors associated with the time until a problematic citation is corrected, we fitted an extended Cox proportional hazards model. This approach models the hazard (instantaneous rate) of correction, where a Hazard Ratio (HR) greater than 1 indicates a faster correction (higher likelihood of correction), whereas an HR less than 1 indicates a slower correction (persistence). For continuous predictors, hazard ratios correspond to a one–standard deviation increase in the log-transformed covariate (log1p, then z-scored). The overall model is statistically significant ($\text{log-likelihood ratio test}:p<0.001$) and demonstrates good predictive accuracy ($concordance=0.684$). We formally assessed multicollinearity using the diagnostic method proposed by Lee and Weissfeld (1996) for the extended Cox proportional hazards model with time-dependent covariates Yul Lee and Weissfeld (1996). This diagnostic was computed from the model-based (non-robust) information matrix, as it is intended to evaluate linear dependence in the covariate design rather than sampling variability of the coefficient estimates. Robust (sandwich) standard errors are reported for inference, but they do not alter the collinearity structure of the covariates. The analysis yielded a maximum Condition Index (CI) of 2.87, well below the recommended threshold of 10. Variance decomposition proportions revealed no coupled loadings across predictors (see Supplementary Table A3), suggesting that the coefficient estimates are not materially affected by multicollinearity. Figure 8 summarizes the hazard ratio estimates, with full model results reported in Appendix Table A4.